### v3

# Saumili Janaa,1, John Kolinskib,2, Detlef Lohsea,c,3, and Vatsal Sanjaye,4

[Uncaptioned image] Impacting spheres: from liquid drops to elastic beads
Saumili Janaa,1, John Kolinskib,2, Detlef Lohsea,c,3, and Vatsal Sanjaye,4
[Uncaptioned image] A liquid drop impacting a non‑wetting rigid substrate spreads laterally, then retracts, and finally jumps off again. An elastic solid, by contrast, undergoes a slight deformation, contacts briefly, and bounces. The impact force on the substrate – crucial for engineering and natural processes – is classically described by Wagner’s (liquids) and Hertz’s (solids) theories. This work bridges these limits by considering a generic viscoelastic medium. Using direct numerical simulations, we study a viscoelastic sphere impacting a rigid, non‑contacting surface and quantify how the elasticity number ($El$, dimensionless elastic modulus) and the Weissenberg number ($Wi$, dimensionless relaxation time) dictate the impact force. We recover the Newtonian liquid response as either $El\to 0$ or $Wi\to 0$, and obtain elastic‑solid behavior in the limit $Wi\to\infty$ and $El\neq 0$. In this elastic‑memory limit, three regimes emerge – capillary‑dominated, Wagner scaling, and Hertz scaling – with a smooth transition from the Wagner to the Hertz regime. Sweeping $Wi$ from 0 to $\infty$ reveals a continuous shift from materials with no memory to materials with permanent memory of deformation, providing an alternate, controlled route from liquid drops to elastic beads. The study unifies liquid and solid impact processes and offers a general framework for the liquid‑to‑elastic transition relevant across systems and applications.
footnotetext: a Physics of Fluids Department, Max Planck Center Twente for Complex Fluid Dynamics, and J. M. Burgers Center for Fluid Dynamics, University of Twente, P.O. Box 217, 7500AE Enschede, the Netherlands.footnotetext: b Institute of Mechanical Engineering, School of Engineering, EPFL, Lausanne, Switzerland.footnotetext: c Max Planck Institute for Dynamics and Self-Organisation, Am Fassberg 17, 37077 Göttingen, Germany.footnotetext: d CoMPhy Lab, Department of Physics, Durham University, Science Laboratories, South Road, Durham DH1 3LE, United Kingdom.footnotetext: 1 s.jana@utwente.nlfootnotetext: 2 john.kolinski@epfl.chfootnotetext: 3 d.lohse@utwente.nlfootnotetext: 4 vatsal.sanjay@comphy-lab.org
1 Introduction
Impacts of spherical bodies on rigid substrates span two classical limits that have long been treated separately: liquid drops 74, 48, 76, 56, 1 and elastic solids 8, 47, 23, 30. Both occur widely in nature and technology, where the normal force on the substrate is often the quantity of interest because it can damage engineered surfaces 3, 12. Drop impact governs processes from inkjet printing 36 and spray cooling/coating 59, 27 to forensics 60, pesticide deposition 22, and soil erosion 43. Impacts of elastic solids arise in hardness testing 71, granular media and suspensions 4, sports 16, and everyday bouncing of soft rubber balls. Despite this breadth, a unifying framework for the impact force across liquid and solid limits has remained elusive.
A falling liquid drop, after impact on a rigid surface deforms and spreads laterally until it reaches its maximum extent. A pronounced peak in the temporal evolution of force occurs at the instance of drop touchdown on the surface due to the inertia of the impact, whereas during droplet spreading this force is much smaller 1. For perfectly wetting surfaces, the liquid sticks to it. However, for non-wetting surfaces, the drop retracts from its maximum spread and generates a Worthington jet which coincides with a second peak in the temporal evolution of force 76, 56. In the inertial regime, Wagner’s theory predicts that the impact force scales as
$F\sim\rho_{l}V_{0}^{2}R_{0}^{2},$ (1)
where $\rho_{l}$ is the density of the media, $V_{0}$ is the impact velocity and $R_{0}$ is the radius of the falling drop 12.
By contrast, the elastic solids undergo slight deformation on impact with the substrate and bounce off following a brief contact with the substrate due to the exerted normal reaction. For such cases, the temporal evolution of force is characterized by a single maximum. Assuming the contact area to be small in comparison to the bead’s size, and considering a non-adhesive contact with small strains within the elastic limit, the situation can be treated as a Hertzian contact problem 30. Thus, Hertz’s theory describes the scaling laws for the impact force in the case of the impact of an elastic bead on a rigid substrate,
$F\sim(GR_{0}^{2})^{2/5}(\rho V_{0}^{2}R_{0}^{2})^{3/5}$ (2)
where $G$ is the modulus of rigidity, $\rho$ is the density of the solid medium, and $R_{0}$ and $V_{0}$ as mentioned before are the radius and impact velocity of the solid elastic bead.
Viscoelastic media – here, soft elastic gels – bridge liquids and solids: when deformed they support both viscous flow and recoverable elastic stress 63. Compared with Newtonian liquids, their rate‑dependent rheology can markedly alter spreading, pinch‑off, and rebound on impact 15, 65. Such soft media are relevant to inkjet printing 9, drop deposition 5, and spray atomization 68. Soft solids such as hydrogels, comprising cross‑linked networks with tunable elasticity, are widely used as biocompatible materials in rapid prototyping 17 and drug delivery 35; see Chen et al. 11 for background.
In this work we parameterize the gel’s elastic response using the elastic modulus $G$ which is the proportionality constant between strain and elastic stresses, and the relaxation time $\lambda$ that sets the decay timescale of those elastic stresses. Upon non-dimensionalizing the governing equations (section 2.1), two material control parameters emerge. The elastocapillary number
$Ec=\frac{GR_{0}}{\gamma}$ (3)
compares the elastic modulus to the Laplace pressure, while the Deborah number
$De=\frac{\lambda}{\sqrt{\rho_{l}R_{0}^{3}/\gamma}}$ (4)
compares the elastic-stress relaxation timescale to the inertio-capillary process time. Here $\gamma$ represents the coefficient of surface tension.
The impact inertia is expressed by the Weber number
$We=\frac{\rho_{l}V_{0}^{2}R_{0}}{\gamma},$ (5)
which compares the inertial and capillary forces. Another important dimensionless control parameter of the system is the Ohnesorge number
$Oh=\frac{\eta_{l}}{\sqrt{\rho_{l}\gamma R_{0}}}$ (6)
which is the ratio of inertio-capillary and the visco-capillary timescales. Here $\eta_{l}$ is the dynamic viscosity of the sphere.
Two combinations of these numbers will be central in what follows. The elasticity number
$El=\frac{Ec}{We}=\frac{G}{\rho_{l}V_{0}^{2}},$ (7)
compares elastic to inertial stresses. The Weissenberg number
$Wi=De\sqrt{We}=\frac{\lambda}{R_{0}/V_{0}},$ (8)
compares the elastic relaxation time $\lambda$ to the impact time $R_{0}/V_{0}$ (see § 2.1).
In this study, we simulate impacts of soft gel spheres on a rigid, non‑contacting substrate using a volume‑of‑fluid, finite‑volume framework. By varying the elastocapillary number ($Ec$) and the Deborah number ($De$), we traverse smoothly from liquid‑like to solid‑like response and compare the resulting force scalings with Wagner’s (liquids) and Hertz’s (elastic solids) theories. We develop an expression for the peak force that transcends the two regimes, using a function of the elasticity parameter to compare the shear modulus with the impact stress. Consistent with these limits, we recover Newtonian‑liquid behavior for $De=0$ or $Ec=0$, while for $De\to\infty$ at sufficiently large $Ec$ the dynamics converge to those of an elastic solid.
2 Numerical Framework
2.1 Problem description and governing equations
We consider an axisymmetric sphere of radius $R_{0}$ approaching a rigid substrate with initial velocity $V_{0}$. For liquid drops the substrate is non‑wetting; for viscoelastic and elastic beads it is non‑contacting. The sphere is a viscoelastic medium of density $\rho_{l}$, dynamic viscosity $\eta_{l}$, elastic modulus $G$, relaxation time $\lambda$, and surface tension coefficient $\gamma$. The surrounding gas has density $\rho_{g}$ and viscosity $\eta_{g}$ (figure 1).
Fig. 1: Schematic: a viscoelastic sphere (radius $R_{0}$) impacts a non‑contacting rigid surface with velocity $V_{0}$. Material properties are $\rho_{l}$, $\eta_{l}$, $G$, $\lambda$, and $\gamma$ for the sphere; $\rho_{g}$ and $\eta_{g}$ for the gas.
Lengths are scaled by $R_{0}$ and time by the inertio‑capillary timescale $\tau_{\gamma}=\sqrt{\rho_{l}R_{0}^{3}/\gamma}$. The corresponding velocity and pressure/stress scales are $u_{\gamma}=\sqrt{\gamma/(\rho_{l}R_{0})}$ and $\sigma_{\gamma}=\gamma/R_{0}$, respectively. Throughout the manuscript, all variables with tilde are non-dimensionalized using the above mentioned scales. The incompressible mass and momentum balances in the viscoelastic phase read
$\bm{\nabla\cdot}\bm{u}=0$ (9)
and,
$\frac{\partial(\rho\bm{u})}{\partial t}+\bm{\nabla\cdot}(\rho\bm{uu})=-\bm{\nabla}p+\bm{\nabla\cdot}(\bm{\sigma_{v}}+\bm{\sigma_{e}})+\bm{f}_{\gamma}$ (10)
where $\bm{f}_{\gamma}$ is the non‑dimensional capillary force density acting at the interface. The Newtonian (viscous) stress is
$\bm{\tilde{\sigma_{v}}}=2Oh\bm{\tilde{\mathcal{D}}}$ (11)
with $\bm{\tilde{\mathcal{D}}}=(\bm{\tilde{\nabla}}\bm{\tilde{u}}+\bm{\tilde{\nabla}}\bm{\tilde{u}}^{T})/2$ denoting the symmetric part of the velocity gradient tensor. Further, the normal force on the substrate is obtained using the rate of change of sphere’s momentum
$\bm{F}(t)=\frac{4}{3}\pi R_{0}^{3}\rho_{l}\frac{dV_{cm}}{dt},$ (12)
where $V_{cm}$ denotes the velocity of the centre of mass of the drop at any instant.
The elastic stresses arise from deformation of the microstructure quantified by the conformation tension $\bm{\mathcal{A}}$ 63. Using the Oldroyd‑B constitutive model 46,
[…]
where the upper‑convected derivative is
$).\stackrel{{\scriptstyle\smash{\raisebox{0.0pt}{$\mkern 8.0mu\bm{\nabla}$}}}}{{\bm{\mathcal{A}}}}\equiv\frac{\partial\bm{\mathcal{A}}}{\partial\tilde{t}}+(\bm{\tilde{u}}\cdot\bm{\tilde{\nabla}})\bm{\mathcal{A}}-2\text{Sym}(\bm{\mathcal{A}}\cdot(\bm{\tilde{\nabla}\tilde{u}}$)). (15)
The Deborah number $De$ quantifies material memory: $De=0$ recovers a Newtonian liquid characterized by $Oh$; $De\to\infty$ yields an elastic solid limit with
$\stackrel{{\scriptstyle\smash{\raisebox{0.0pt}{$\mkern 8.0mu\bm{\nabla}$}}}}{{\bm{\mathcal{A}}}}=0$ (16)
and Oldroyd‑B equivalent to a neo‑Hookean solid 63.
Despite being widespread due to simplicity, the Oldroyd-B model suffers from certain limitations 63, 18. It fails to capture the shear-thinning behaviour in viscoelastic fluids completely, and erroneously predicts unbounded stress growth in strong extensional flows 40. This limitation can be addressed by incorporating the finite polymer extension (FENE-P model)6. Also various other extensions of the Oldroyd-B model have been developed 66 like the Phan-Thien-Tanner (PTT)67 model to account for such non-linearities. However, since we are not dealing with strong extensional flows, we restrict ourselves to the Oldroyd-B model in this study, as it is sufficient to describe our case.
2.2 Numerical methods and simulations
We solve the above equations with Basilisk C 50, using a one‑fluid formulation with surface tension as a singular interfacial force 72, 7. The liquid–gas interface is tracked by a volume‑of‑fluid (VoF) color function $\psi$ advected by
$\frac{\partial\psi}{\partial t}+\bm{\nabla}\cdot(\psi,\bm{u})=0,$ (17)
with $\psi=1$ in liquid, $\psi=0$ in gas, and $0<\psi<1$ at the interface. Mixture properties are
$\rho=\psi\rho_{l}+(1-\psi)\rho_{g},$ (18)
$\eta=\psi\eta_{l}+(1-\psi)\eta_{g},$ (19)
with fixed ratios $\rho_{r}=\rho_{g}/\rho_{l}=10^{-3}$ and $\eta_{r}=\eta_{g}/\eta_{l}=10^{-2}$ and a small sphere Ohnesorge number, $Oh=10^{-2}$. A geometric VoF reconstruction applies capillary forces as
$\bm{f}_{\gamma}\approx\gamma\kappa\bm{\nabla}\psi,$ (20)
where the curvature $\kappa$ is computed via height functions 51. Explicit surface‑tension forcing imposes the standard capillary time‑step constraint 52; the explicit update of $\bm{\sigma_{e}}$ adds a typically milder constraint.
At the substrate we impose no‑penetration and no‑slip, and a zero normal pressure gradient. To enforce a non‑contacting (superhydrophobic) condition we set $\psi=0$ at the wall, maintaining a thin air cushion 28, 57. For liquid drops, enforcing this air cushion results in a non-wetting (superhydrophobic) substrate, while for elastic spheres this results in a non-contacting substrate. We stress that contact initiation in soft–solid impacts is generically air‑mediated and can proceed annularly or patchily with a non‑monotonic initial contact radius; see Zheng et al. (2021)78 for direct observations of air-mediated contact in compliant‑hemisphere impacts. Top and lateral boundaries use outflow (ambient pressure, zero tangential stress, zero normal velocity gradient). Boundaries are positioned far enough to avoid spurious confinement effects. The axisymmetric domain size is $8R_{0}\times 8R_{0}$. We employ quadtree adaptive mesh refinement (AMR) 52, 53 with maximal refinement at the interface and in regions of large velocity gradients. Wavelet‑based error control uses tolerances $10^{-3}$ for $\bm{u}$, $\psi$, $\kappa$, and $\bm{\mathcal{A}}$. Grid‑independence tests confirm convergence. Unless stated otherwise, the minimum cell size is $\Delta=R_{0}/512$ (i.e. 512 cells per radius on a uniform equivalent grid), increased to $\Delta=R_{0}/2048$ when required. Further numerical details can be found in Popinet 53, Sanjay 57 and Dixit et al. 18.
3 Wagner versus Hertz: Permanent‑Memory Impacts
In this section, we quantify the solid‑impact limit by taking $De\to\infty$, so the material retains its deformation memory over the process time. In the numerics, we keep a small background viscosity, so the spheres are Kelvin–Voigt solids rather than perfectly elastic; this facilitates comparison with inertial liquid impacts at finite $Oh$ and avoids the numerical breakdown of the inviscid (Euler)–elastic limit. We therefore approach the purely elastic response by letting $Oh\to 0$.
We sweep the $(Ec,We)$ space over $We\in[1,10^{3}]$ and $Ec\in[10^{-1},10^{4}]$. The normal reaction on the substrate is computed from the drop’s momentum balance (eq. 12). For liquid drops on non-wetting substrates, $F(t)$ exhibits two peaks: an inertial peak at touchdown and a later peak associated with the formation of a Worthington jet 76. The second, jetting peak occurs at a time $t_{2}\gg t_{\max}$, with the ratio $t_{2}/t_{\max}\sim\sqrt{We}$ 56, reflecting the phase difference between recoil and jet formation. In contrast, for elastic spheres the loading and unloading remain nearly in phase: the entire contact–rebound cycle fits within $t\lesssim 2t_{\max}$ and $F(t)$ displays only a single, almost symmetric peak in this interval. To compare liquids and solids and to track the transition, we therefore focus on the first (inertial) peak $F_{\max}$, non-dimensionalized as $F_{\max}/(\rho_{l}V_{0}^{2}R_{0}^{2})$; the corresponding time is $t_{\max}$.
Fig. 2: (a) Phase space in the EcEc-WeWe plane illustrating the range of simulations conducted in this work colored according to the elasticity number $El=Ec/We$. The four highlighted symbols locate typical cases representing the range of parameters explored. We chose ($We,Ec$) == (b) ($5,1$), (c) ($500,2$), (d) ($5,1000$), (e) ($500,1000$). For each case, the color scheme of each snapshot represents the magnitude of the velocity normalized by the impact velocity, alongside the corresponding force history $F(t)/(\rho_{l}V_{0}^{2}R_{0}^{2})$ plotted versus $t/t_{\max}$ (right). The force traces are plotted up to $t/t_{\max}=2$: for the liquid-drop reference, the second peak associated with the Worthington jet 76, 56 occurs at later times $t_{2}\gg t_{\max}$ (with $t_{2}/t_{\max}\sim\sqrt{We}$) and is therefore outside the plotted window.
Figure 2 shows representative cases across the parameter space. At low $Ec$ the sphere flows and behaves liquid‑like (figs. 2b,c; $Ec=1,2$). At low $Ec$ and low $We$, capillarity is significant (fig. 2b). At high $Ec$ the sphere deforms slightly and rebounds after a short contact (figs. 2d; $Ec=1000$). Increasing $We$ at fixed $Ec$ effectively softens the response (reduces $El$) and increases the contact duration (figs. 2d). The force traces reflect this evolution: for large $El$ (high $Ec$, low $We$) $F(t)$ is nearly symmetric, as in elastic impacts, while decreasing $El$ (e.g. by increasing $We$) skews $F(t)$ in the manner typical of liquid impacts 76, 56. The peak magnitude also varies appreciably across cases. Since force and time are related as $F_{max}/(\rho_{l}V_{0}^{2}R_{0}^{2})\cdot t_{max}/(R_{0}/V_{0})\sim 1$ for any impacting sphere in general, $t_{max}$ reduces accordingly with an increase in $F_{max}$.
The dependence of $F_{\max}$ on $We$ and $Ec$ is summarized in figure 3. For $Ec\lesssim 1$, $F_{\max}$ follows the liquid‑impact trend with the low‑$We$ correction 76,
$\frac{F_{\max}}{\rho_{l}V_{0}^{2}R_{0}^{2}}\approx\frac{3.2}{We}+3.24.$ (21)
As $Ec$ increases, $F_{\max}$ rises, most strongly at low $We$. At sufficiently large $Ec$, $F_{\max}/(\rho_{l}V_{0}^{2}R_{0}^{2})$ decreases with $We$ with a log–log slope $\simeq-2/5$, indicating a transition from eq. (21) to $F_{\max}/(\rho_{l}V_{0}^{2}R_{0}^{2})\sim We^{-2/5}$. At fixed $We$ (figure 3b), $F_{\max}$ is nearly constant at small $Ec$ and then increases steadily with $Ec$, with higher magnitudes at lower $We$.
Fig. 3: Peak force in the elastic-memory limit ($De\to\infty$): (a) Variation of the normalized peak force, $F_{\max}/(\rho_{l}V_{0}^{2}R_{0}^{2})$, with the Weber number $We=\rho_{l}V_{0}^{2}R_{0}/\gamma$ for different elastocapillary numbers $Ec=GR_{0}/\gamma$. For $Ec\lesssim\mathcal{O}(1)$ the data follow the liquid-impact result: a high-WeWe Wagner plateau $\simeq 3.24$, with the low-WeWe correction $F_{\max}/(\rho_{l}V_{0}^{2}R_{0}^{2})\approx 3.2/We+3.24$ (dashed line, eq. (21)). As $Ec$ increases, $F_{\max}$ rises, most clearly at low $We$, and for sufficiently large $Ec$ the curves acquire a log–log slope 2/5-2/5, i.e. $F_{\max}/(\rho_{l}V_{0}^{2}R_{0}^{2})\sim We^{-2/5}$ at fixed EcEc, consistent with the approach to Hertzian elastic contact. (b) Dependence on EcEc at fixed WeWe (curves labelled by WeWe). At small EcEc all series collapse to the liquid-like level ($\approx 3.24$); above a WeWe-dependent crossover, $F_{\max}$ increases monotonically with $Ec$, following $F_{\max}/(\rho_{l}V_{0}^{2}R_{0}^{2})\sim Ec^{2/5}$, again, consistent with the approach to Hertzian elastic contact. Together, (a,b) show a continuous evolution from Wagner (liquid) to Hertz (elastic) behavior as EcEc increases.
Collapsing the data using $El$ (figure 4a) reveals two regimes. For $El\lesssim 1$ and sufficiently large $We$, the data lie near Wagner’s constant level, $F_{\max}/(\rho_{l}V_{0}^{2}R_{0}^{2})\approx 3.24$. At small $We$, inertia competes with capillarity and the low‑$We$ correction in eq. (21) is required 57. For $El\gtrsim 1$, all points collapse onto a single master curve with slope $2/5$:
$\frac{F_{\max}}{\rho_{l}V_{0}^{2}R_{0}^{2}}\sim El^{2/5}\sim\left(\frac{G}{\rho_{l}V_{0}^{2}}\right)^{2/5},$ (22)
consistent with Hertz scaling for elastic impacts. Notably, surface tension does not enter this high‑$El$ law, as expected for elastic solids. The transition from the Wagner (liquid) to the Hertz (elastic) regime is smooth. A contour map over $(We,Ec)$ (figure 4b) visualizes the continuous variation of $F_{\max}$ across the space.
Fig. 4: Unified scaling and regime map: (a) Collapse of the normalized peak force versus the elasticity number $El=Ec/We=G/(\rho_{l}V_{0}^{2})$. For $El\lesssim 1$ and sufficiently large $We$ the data sit on the Wagner plateau $F_{\max}/(\rho_{l}V_{0}^{2}R_{0}^{2})\approx 3.24$; deviations at very small WeWe reflect capillary corrections in eq. (21). For $El\gtrsim 1$ all cases collapse onto a single power law with slope $2/5$, $F_{\max}/(\rho_{l}V_{0}^{2}R_{0}^{2})\sim El^{2/5}$ (eq. (22)), the hallmark of Hertz scaling. (b) Contours of $F_{\max}/(\rho_{l}V_{0}^{2}R_{0}^{2})$ in the $(We,Ec)$ plane (symbols: simulation points). The dashed guide $El=1$ marks the smooth crossover from the Wagner region (lower right) to the Hertz region (upper left); the low‑$We$ corner is capillary‑dominated and requires the correction in eq. (21). The map visualizes the continuous transition from liquid‑like to solid‑like impact forces as WeWe and EcEc are varied.
4 Theory
In this section we identify the asymptotic limits of the peak impact force and develop a unified predictive model. First, we derive scaling expressions for $F_{\max}$ in the two extreme regimes – an elastic contact limit versus a hydrodynamic impact limit – and then combine these results to propose a single predictive expression for the dimensionless maximum impact force.
4.1 Purely Elastic Limit: Hertz Contact Theory
Consider a solid elastic sphere of radius $R_{0}$, mass $m$, and elastic modulus $G$ (shear modulus, assuming an incompressible material) impacting a rigid flat surface with speed $V_{0}$. Upon contact, the sphere deforms and a normal force $F$ develops according to Hertz’s contact law. For a sphere indenting a half-space, the force–indentation relation is given by the 3/2-power law of classical Hertz contact mechanics 8,
$F(\delta)=m\ddot{\delta}=-\frac{4}{3}E^{*}\sqrt{R_{0}}\delta^{3/2},$ (23)
where $\delta(t)$ is the indentation depth and $E^{*}$ is the effective Young’s modulus of the contacting pair. For a sphere against a rigid flat, $E^{*}=2G/(1-\nu)$; taking Poisson’s ratio $\nu\approx 0.5$ for an incompressible solid, we get $E^{*}\approx 4G$, so the prefactor $(4/3)E^{*}\sqrt{R_{0}}$ in eq. (23) is about $(16/3)G\sqrt{R_{0}}$.
In the ideal elastic limit (no dissipation), the sphere will momentarily come to rest at maximum compression, converting all its kinetic energy into elastic deformation energy. Using energy conservation between the moment of impact and the instant of maximum indentation $\delta_{\max}$ (when $\dot{\delta}=0$), we have:
$\frac{1}{2}mV_{0}^{2}=\int_{0}^{\delta_{\max}}F(\delta)\,d\delta\,.$ (24)
Substituting the Hertz law for $F(\delta)$ and performing the integration yields the elastic energy stored at indentation $\delta_{\max}$:
$\frac{1}{2}mV_{0}^{2}$ $=\int_{0}^{\delta_{\max}}\frac{16}{3}G\sqrt{R_{0}}\delta^{3/2}d\delta$ (25)
$=\frac{16}{3}G\sqrt{R_{0}}\cdot\frac{2}{5}\delta_{\max}^{5/2}$ (26)
$=\frac{32}{15}G\sqrt{R_{0}}\delta_{\max}^{5/2}.$ (27)
Rearranging this result to solve for the peak indentation $\delta_{\max}$ gives:
$\delta_{\max}=\left(\frac{15\,mV_{0}^{2}}{64\,G\sqrt{R_{0}}}\right)^{2/5}.$ (28)
The maximum force occurs at $\delta=\delta_{\max}$. Substituting the expression for $\delta_{\max}$ in eq. 23, we get
$F_{\max}=\frac{16}{3}G\sqrt{R_{0}}\left(\frac{15\,mV_{0}^{2}}{64\,G\sqrt{R_{0}}}\right)^{3/5}.$ (29)
Normalizing with the inertial force scale $\rho_{l}V_{0}^{2}R_{0}^{2}$, we get
$\frac{F_{\max}}{\rho_{l}V_{0}^{2}R_{0}^{2}}=\frac{16}{3}\left(\frac{5\pi}{16}\right)^{3/5}\left(\frac{G}{\rho_{l}V_{0}^{2}}\right)^{2/5}\approx 5.3\,El^{2/5},$ (30)
which is consistent with our large $El$ results, cf. figure 4a.
4.2 Purely Liquid Limit: Wagner Impact Theory
At the opposite extreme ($Ec\to 0$) the sphere behaves as a liquid drop, and its impact dynamics are governed by inertia and capillarity in the classical Wagner limit 12, 76. Instead of an elastic compression, the drop undergoes rapid localized deformation at the moment of impact: the south pole flattens against the substrate while the remainder of the drop (including the north pole) is still moving downward at nearly the impact speed. The vertical momentum of the drop’s center of mass is redirected into a radial outflow along the substrate, causing a small “wetted” area to grow outward from the impact point56. This Wagner-type mechanism – a thin spreading lamella initiated at the contact point 20, 26, 45, 44 – contrasts sharply with the distributed Hertzian contact of an elastic solid. It produces a pronounced impulsive force at touchdown, as the drop’s momentum is arrested over a short time and small area. The first force peak thus originates from pure inertial impingement of the liquid on the surface1. We have analyzed this case in past, for details see the references Zhang et al. 76, Sanjay et al. 56, Sanjay and Lohse 1.
During the very early stage ($t\sim\tau_{\rho}=D_{0}/V_{0}$), the normal force rises sharply to its first maximum $F_{\max}$ as the drop’s inertia is transferred to the substrate. At this moment the deformation is still localized: the contact radius has grown only to the order of the drop’s initial radius. In fact, experiments confirm that at the peak force time $t_{\max}$, the spread diameter $D_{f}(t_{\max})$ is approximately equal to the initial drop diameter $D_{0}$, consistent with early-time self-similarity of the impact 56, 1, 38, Philippi_Lagrée_Antkowiak_2016. Wagner’s inviscid theory predicts that the peak force scales with the inertial pressure on the drop’s footprint. Non-dimensionalizing $F_{\text{max}}$ by $\rho_{l}V_{0}^{2}R_{0}^{2}$ (with $\rho_{l}$ the liquid density) yields a constant of order unity. Indeed, for large Weber numbers (negligible surface tension), simulations and experiments find $F_{\text{max}}/(\rho_{l}V_{0}^{2}R_{0}^{2})\approx 3.24$, cf. figure 3a of Zhang et al. 76. At lower $We$, surface tension enhances the impact, and the peak force is increases with decreasing $We$ (following a $F_{\max}/(\rho_{l}V_{0}^{2}R_{0}^{2})\sim We^{-1}$ correction in this regime). This initial peak is inertia-dominated and is relatively insensitive to liquid viscosity: $F_{\text{max}}$ remains nearly constant for drops with viscosity up to about $100\times$ that of water 56. Since our spheres have a very low background viscosity ($Oh=10^{-2}$), in the limit $Ec\to 0$ our results for $F_{\text{max}}$ follow the same trends. Only for highly viscous drops (Ohnesorge number $Oh\gtrsim 1$) does viscous dissipation significantly attenuate the first peak, reflecting the fact that most of the drop’s momentum is redirected (and the force generated) before substantial viscous effects have time to act, cf. Sanjay and Lohse 1.
4.3 Predictive Interpolating Model for Maximum Impact Force
For intermediate conditions ($El\sim\mathcal{O}\left(1\right)$), the sphere’s deformation and the fluid’s inertia both contribute, and the peak force deviates from either pure Hertz or Wagner scaling alone. We therefore express the dimensionless peak force as a weighted transition between the two asymptotic contributions following the approach of Sanjay & Lohse (2025)1,
$\frac{F_{\max}}{\rho_{l}V_{0}^{2}R_{0}^{2}}=5.3\,\text{f}\left(El\right)\,El^{2/5}+(1-\text{f}\left(El\right))\Big(\frac{3.2}{We}+3.24\Big),$ (31)
with a smooth transition function $\text{f}\left(El\right)$ based on tanh function defined as
$\text{f}(El)=\frac{1+\text{tanh}(\frac{El-a}{b})}{2},$ (32)
where the $We$-dependant parameters $a\approx 1.3$ for $We\to 1$ and $\approx 0.5$ for $We\gtrsim 10$ and $b\approx 1.4$ for $We\to 1$ and $\approx 0.2$ for $We\gtrsim 10$ (see 25 for details of this fit). Here, $a$ measures the critical $El$ at transition from Wagner’s to Hertz’s scaling (fig.4). The width of this transition is indicated by $b$. The remaining coefficients in eq. (31) are fixed a priori: the Hertz prefactor 5.3 follows directly from the elastic analysis above (without fitting), while the constants 3.24 (Wagner plateau) and 3.2 (the $We^{-1}$ capillary correction) are taken from the Newtonian impact model of Sanjay & Lohse (2025)1. This construction ensures a continuous interpolation between the Hertz and Wagner limits. In spirit it follows the additive scaling approach of Sanjay & Lohse (2025)1 for drop impacts, but unlike their model – which includes separate viscous regimes – here only the two primary regimes (elastic vs. inertial) are needed. The resulting formula smoothly bridges the two asymptotes and correctly reproduces the peak-force scaling in both limits (this predictive curve is plotted in fig. 4a for comparison).
5 Influence of Elastic Stress Relaxation
Fig. 5: Relaxing material memory at fixed WeWe and $El$. Evolution of shape (left) and force (right) when WiWi decreases from $\infty$ to 0 at $We=100$ and $El=40$: (a) $Wi\to\infty$ (elastic‑memory limit); (b) $Wi=10^{-1}$; (c) $Wi=10^{-2}$; (d) $Wi=0$ (Newtonian). For each case, the color scheme of each snapshot represents the magnitude of the velocity normalized by the impact velocity, alongside the corresponding force history $F(t)/(\rho_{l}V_{0}^{2}R_{0}^{2})$ plotted versus $t/t_{\max}$ (right). As WiWi decreases, contact time increases and $F(t)$ becomes increasingly liquid‑like with a reduced $F_{\max}$.
Fig. 6: Peak force versus elasticity number at different WiWi (all at $We=100$). The black horizontal line indicates the Wagner plateau ($\simeq 3.24$); the dashed guide has slope $2/5$ (Hertz). Increasing WiWi shifts the departure from the plateau to lower ElEl and drives the curves toward the Hertz master law; at $Wi=0$ all data sit on the Wagner level.
The results in § 3 established the two asymptotic force scalings for impacts with permanent memory ($De\to\infty$): a Wagner plateau at small elasticity number and a Hertz law at large elasticity number (figure 4). These two limits also bound impacts when the material memory is finite: relaxing the memory shifts the response continuously from solid‑like to liquid‑like, with the transition controlled by the non‑dimensional relaxation time. We quantify the memory effect with the Weissenberg number, $Wi$ (eq. 8), which compares the elastic relaxation time $\lambda$ to the impact time $R_{0}/V_{0}$ (see § 2.1). While $De$ compares $\lambda$ to the inertio‑capillary time, $Wi$ is the more natural process‑time measure here and increases with the degree to which elastic stresses persist during impact 49. Thus $Wi=0$ ($De=0$) recovers a Newtonian liquid, whereas $Wi\to\infty$ ($De\to\infty$) yields an elastic‑memory limit.
Figure 5 (fixed $We=100$ and $El=40$) visualizes the progressive loss of elastic behavior as $Wi$ decreases from $Wi\to\infty$ to 0. In the large‑$Wi$ limit, the bead contacts briefly and rebounds; the force trace is nearly symmetric with a large peak, characteristic of Hertz‑like loading. Reducing $Wi$ increases the contact time and skews $F(t)$ towards a liquid‑like evolution with a much smaller peak. At $Wi=0$ the material has no memory and behaves as a Newtonian liquid: the sphere spreads and flows, and $F(t)$ exhibits the familiar asymmetric shape.
To quantify the role of memory, we plot $F_{\max}/(\rho_{l}V_{0}^{2}R_{0}^{2})$ versus $El=Ec/We=G/(\rho_{l}V_{0}^{2})$ for several $Wi$ at $We=100$ in fig. 6. At $Wi=0$, the data follows the Newtonian liquid level established in § 3. On the other hand, as $Wi\to\infty$, the points converge to the elastic‑memory master curve of § 3, namely the Hertz scaling $F_{\max}/(\rho_{l}V_{0}^{2}R_{0}^{2})\approx 5.3El^{2/5}$ at large $El$. Between these limits, weaker memory (smaller $Wi$) delays the transition to Hertz scaling, while stronger memory (larger $Wi$) makes the elastic response apparent already for softer spheres. Notably, even at $Wi\to\infty$ the plateau persists for $El\ll 1$, because the modulus is introduced exclusively through $El$: vanishing $G$ implies a liquid‑like bound irrespective of memory.
Taken together with § 3, these results show that both material stiffness (via $El$) and material memory (via $Wi$) govern the peak impact force: the Wagner and Hertz laws remain the bounding asymptotes, while $Wi$ sets how rapidly the system transitions between them.
Conclusions & outlook
In this work, we investigate the impact of a viscoelastic sphere on a non-contacting rigid surface and chart a continuous transition in impact dynamics from liquid-like to solid-like behavior by tuning the material parameters. The primary peak force $F_{\max}$ (associated with the inertial impact) smoothly crosses over from Wagner’s inertial-drop scaling to Hertz’s elastic-contact scaling as the elasticity number $El=Ec/We$ increases. For small $El$ (liquid-like response), we reproduce $F_{\max}/(\rho_{l}V_{0}^{2}R_{0}^{2})\to 3.24$ matching the constant plateau from Wagner’s theory. In contrast, for large $El$ (elastic-dominated regime), $F_{\max}/(\rho_{l}V_{0}^{2}R_{0}^{2})$ grows following a power-law $\approx 5.3El^{2/5}$, consistent with Hertz’s prediction for elastic spheres. These two limiting behaviors bound the force response, and the transition between them is gradual rather than abrupt. The Weissenberg number $Wi=De\,\sqrt{We}$, which quantifies the polymer’s relaxation time relative to the impact time, governs this memory-driven crossover: as $Wi$ increases from 0 (no elastic memory) to $\infty$ (permanent memory), the peak-force scaling shifts continuously from the Wagner limit to the Hertz limit. Thus, by adjusting $Wi$, one can smoothly interpolate between liquid-drop and elastic-solid impact outcomes.
The modeling choices in this work were made to isolate the physics of the liquid–to–elastic transition and enable direct comparison to the classical limits. The viscoelastic sphere obeys Oldroyd‑B, which neglects finite microstructure extensibility and shear‑thinning; this is acceptable for our flow history but can be systematically relaxed with constitutive models that incorporate finite extensibility and rate‑dependent viscosity 63, 40, 67, 66. The substrate is non‑contacting, so the thin gas layer is present but not explicitly resolved with lubrication and wetting dynamics; prior work shows that air cushioning, skating on a gas film, and nanoscale first contact depend sensitively on slip, compressibility and rarefaction 61, 70, 69, 29, 73, 19, 33, 14, 21, 31, 77, 64. We stress that in the present work, the substrate is assumed to remain perfectly rigid. Small elastic deformations of the substrate can, however, modify the intervening gas layer 32, which in turn may alter the impact dynamics and the measured force signature of the impacting sphere.Ma Furthermore, the onset of contact will modify the shear stress at the interface - including effects such as adhesion - that can modify the peak force, and will alter the stresses upon rebound. Finally, a small background viscosity renders the ($De\to\infty$) limit to be of Kelvin–Voigt type rather than ideally elastic; classical analyses quantify how viscoelastic dissipation and elastic waves perturb Hertzian impact 47, 23. These assumptions do not alter the governing exponents (Wagner versus Hertz) but they may affect quantitative prefactors and very‑short‑time, near‑contact details.
Several direct extensions can sharpen and generalize these results. Experiments with soft hydrogel/elastomer beads at moderate $V_{0}$ can probe the crossover regime and test the full $F(t)$ waveform (peak magnitude, rise time, symmetry/skewness), leveraging recent studies on soft‑solid and gel impacts and elastohydrodynamic bouncing 42, 24, 37, 39, 13, 2, 30, and using established force‑measurement protocols from liquid‑drop impacts 41, 34, 75, 76, 56. Furthermore, incorporating a thin‑gas lubrication model with dynamic wetting (including slip and, if needed, rarefaction/compressibility) will resolve when and how contact initiates and how this feeds back on the very‑early‑time force 78, 61, 29, 73, 33, 19, 38, 14, 21, 64. Linking impact memory with lubricated‑impact and wetting transitions will place the near‑contact force history on firmer ground 62, 54, 58, 55, 10.
Author contributions
V.S., J.K., and D.L. conceived the study. S.J. and V.S. planned the numerical simulations. S.J. performed the simulations and analyzed the data. V.S. and D.L. developed the theory. S.J. and V.S. designed the structure of the manuscript. S.J., V.S., D.L. and J.K. wrote the manuscript. V.S. and D.L. supervised the project. All authors discussed the results and approved the final manuscript.
Conflicts of interest
There is no conflict of interest.
Data availability
All codes used in this work is available as an open-source repository at https://github.com/comphy-lab/Soft-Sphere-Impacts 25.
Acknowledgements
We would like to thank Vincent Bertin, Pierre Chantelot, Maziyar Jalaal, Andrea Prosperetti, and Jacco Snoeijer for discussions. This work was carried out on the national e-infrastructure of SURFsara, a subsidiary of SURF cooperation, the collaborative ICT organization for Dutch education and research. This work was sponsored by NWO - Domain Science for the use of supercomputer facilities. This work was supported by NWO-Canon grant FIP-II grant. V.S. acknowledges start-up funding from Durham University.
Notes and references
  • V. Sanjay and D. Lohse (2025) Unifying theory of scaling in drop impact: forces and maximum spreading diameter. Phys. Rev. Lett. 134, pp. 104003. External Links: Document Cited by: §1, §1, §4.2, §4.2, §4.3, §4.3.
  • E. A. Agüero, L. F. L. Alventosa, D. M. Harris, and C. A. Galeano-Rios (2022) Impact of a rigid sphere onto an elastic membrane. Proc. R. Soc. Lond. A 478 (2266), pp. 20220340. Cited by: Conclusions & outlook.
  • M. Ahmad, M. Schatz, and M. Casey (2013) Experimental investigation of droplet size influence on low pressure steam turbine blade erosion. Wear 303, pp. 83–86. External Links: Document Cited by: §1.
  • B. Andreotti, Y. Forterre, and O. Pouliquen (2013) Granular media: between fluid and solid. Cambridge University Press. Cited by: §1.
  • V. B. Bergeron, D. B. Bonn, J. M. Martin, and V. L. Louis (2000) Controlling droplet deposition with polymer additives. Nature 405, pp. 772–775. External Links: Document Cited by: §1.
  • R.B. Bird, P.J. Dotson, and N.L. Johnson (1980) Polymer solution rheology based on a finitely extensible bead—spring chain model. J. Non-Newtonian Fluid Mech. 7 (2), pp. 213–235. External Links: ISSN 0377-0257, Document Cited by: §2.1.
  • J. U. Brackbill, D. B. Kothe, and C. Zemach (1992) A continuum method for modeling surface tension. J. Comput. Phys. 100 (2), pp. 335–354. Cited by: §2.2.
  • H. Hertz (1882) Ueber die berührung fester elastischer körper.. Journal für die reine und angewandte Mathematik 1882, pp. 156–171 (en). External Links: Link, Document, ISSN 0075-4102, 1435-5345 Cited by: §1, §4.1.
  • P. Calvert (2001) Inkjet printing for materials and devices. Chem. Mater. 13, pp. 3299–3305. External Links: Document Cited by: §1.
  • P. Chantelot and D. Lohse (2021) Drop impact on superheated surfaces: short-time dynamics and transition to contact. J. Fluid Mech. 928, pp. A36. Cited by: Conclusions & outlook.
  • D. Chen, Q. Wen, P. Janmey, J. Crocker, and A. Yodh (2010) Rheology of soft materials. Annu. Rev. Condens. Matter Phys 1, pp. 301–22. External Links: Document Cited by: §1.
  • X. Cheng, T.-P. Sun, and L. Gordillo (2022) Drop impact dynamics: impact force and stress distributions. Annu. Rev. Fluid Mech. 54, pp. 57–81. Cited by: §1, §1, §4.2.
  • F. Chevy, A. Chepelianskii, D. Quéré, and E. Raphaël (2012) Liquid Hertz contact: softness of weakly deformed drops on non-wetting substrates. Europhys. Lett. 100 (5), pp. 54002. Cited by: Conclusions & outlook.
  • M. V. Chubynsky, K. I. Belousov, D. A. Lockerby, and J. E. Sprittles (2020) Bouncing off the walls: the influence of gas-kinetic and van der Waals effects in drop impact. Phys. Rev. Lett. 124, pp. 084501. Cited by: Conclusions & outlook, Conclusions & outlook.
  • C. Clasen, J. Eggers, M. A. Fontelos, J. Li, and G. H. Mckinley (2006) The beads-on-string structure of viscoelastic threads. J. Fluid Mech. 556, pp. 283–308. External Links: Document Cited by: §1.
  • R. Cross (1999) The bounce of a ball. Am. J. Phys. 67 (3), pp. 222–227. External Links: Document, ISSN 0002-9505 Cited by: §1.
  • B. Derby (2012) Printing and prototyping of tissues and scaffolds. Science 338, pp. 921–6. External Links: Document Cited by: §1.
  • A. K. Dixit, A. Oratis, K. Zinelis, D. Lohse, and V. Sanjay (2025) Viscoelastic Worthington jets and droplets produced by bursting bubbles. J. Fluid Mech. 1010, pp. A2. External Links: ISSN 0022-1120, 1469-7645, Document Cited by: §2.1, §2.2.
  • M. M. Driscoll and S. R. Nagel (2011) Ultrafast interference imaging of air in splashing dynamics. Phys. Rev. Lett. 107 (15), pp. 154502. Cited by: Conclusions & outlook, Conclusions & outlook.
  • J. Eggers, M. A. Fontelos, C. Josserand, and S. Zaleski (2010) Drop dynamics after impact on a solid wall: theory and simulations. Phys. Fluids 22 (6), pp. 062101. Cited by: §4.2.
  • P. García-Geijo, G. Riboux, and J. Gordillo (2024) The skating of drops impacting over gas or vapour layers. J. Fluid Mech. 980, pp. A35. Cited by: Conclusions & outlook, Conclusions & outlook.
  • L. He, L. Ding, B. Li, W. Mu, P. Li, and F. Liu (2021) Optimization strategy to inhibit droplets rebound on Pathogen-Modified hydrophobic surfaces. ACS Appl. Mater. Interfaces 13 (32), pp. 38018–38028. Cited by: §1.
  • S. C. Hunter (1957) Energy absorbed by elastic waves during impact. J. Mech. Phys. Solids 5 (3), pp. 162–171. External Links: ISSN 0022-5096, Document Cited by: §1, Conclusions & outlook.
  • M. Jalaal, C. Seyfert, B. Stoeber, and N. J. Balmforth (2018) Gel-controlled droplet spreading. J. Fluid Mech. 837, pp. 115–128. Cited by: Conclusions & outlook.
  • S. Jana and V. Sanjay (2026) Code repository: Soft Sphere Impacts. Note: https://github.com/comphy-lab/Soft-Sphere-Impacts Cited by: §4.3, Data availability.
  • C. Josserand and S. T. Thoroddsen (2016) Drop impact on a solid surface. Annu. Rev. Fluid Mech. 48, pp. 365–391. Cited by: §4.2.
  • J. Kim (2007) Spray cooling heat transfer: the state of the art. Int. J. Heat Fluid Flow 28 (4), pp. 753–767. External Links: Document, ISSN 0142-727X Cited by: §1.
  • J. M. Kolinski, L. Mahadevan, and S. M. Rubinstein (2014) Drops can bounce from perfectly hydrophilic surfaces. Europhys. Lett. 108, pp. 24001. Cited by: §2.2.
  • J. M. Kolinski, S. M. Rubinstein, S. Mandre, M. P. Brenner, D. A. Weitz, and L. Mahadevan (2012) Skating on a film of air: drops impacting on a surface. Phys. Rev. Lett. 108, pp. 074503. Cited by: Conclusions & outlook, Conclusions & outlook.
  • V. Bertin (2024) Similarity solutions in elastohydrodynamic bouncing. J. Fluid Mech. 986, pp. A13. External Links: Link, Document Cited by: §1, §1, Conclusions & outlook.
  • K. Langley, E. Q. Li, and S. T. Thoroddsen (2017) Impact of ultra-viscous drops: air-film gliding and extreme wetting. J. Fluid Mech. 813, pp. 647–666. Cited by: Conclusions & outlook.
  • K. R. Langley, A. A. Castrejón-Pita, and S. T. Thoroddsen (2020) Droplet impacts onto soft solids entrap more air. Soft Matter 16, pp. 5702–5710. External Links: Document Cited by: Conclusions & outlook.
  • E. Q. Li, I. U. Vakarelski, and S. T. Thoroddsen (2015) Probing the nanoscale: the first contact of an impacting drop. J. Fluid Mech. 785, pp. R2. Cited by: Conclusions & outlook, Conclusions & outlook.
  • J. Li, B. Zhang, P. Guo, and Q. Lv (2014) Impact force of a low speed water droplet colliding on a solid surface. J. Appl. Phys. 116 (21), pp. 214903. Cited by: Conclusions & outlook.
  • J. Li and D. Mooney (2016) Designing hydrogels for controlled drug delivery. Nature Rev. Mater. 1, pp. 16071. External Links: Document Cited by: §1.
  • D. Lohse (2022) Fundamental fluid dynamics challenges in inkjet printing. Annu. Rev. Fluid Mech. 54, pp. 349–382. Cited by: §1.
  • L.-H. Luu and Y. Forterre (2009) Drop impact of yield-stress fluids. J. Fluid Mech. 632, pp. 301–327. Cited by: Conclusions & outlook.
  • S. Mandre, M. Mani, and M. P. Brenner (2009) Precursors to splashing of liquid droplets on a solid surface. Phys. Rev. Lett. 102, pp. 134502. External Links: Document, Link Cited by: §4.2, Conclusions & outlook.
  • G. Martouzet, L. Jørgensen, Y. Pelet, A.-L. Biance, and C. Barentin (2021) Dynamic arrest during the spreading of a yield stress fluid drop. Phys. Rev. Fluids 6 (4), pp. 044006. Cited by: Conclusions & outlook.
  • G. H. McKinley and T. Sridhar (2002) Filament-stretching rheometry of complex fluids. Annu. Rev. Fluid Mech. 34, pp. 375–415. External Links: Document, ISSN 1545-4479 Cited by: §2.1, Conclusions & outlook.
  • B. R. Mitchell, J. C. Klewicki, Y. P. Korkolis, and B. L. Kinsey (2019) The transient force profile of low-speed droplet impact: measurements and model. J. Fluid Mech. 867, pp. 300–322. Cited by: Conclusions & outlook.
  • S. Mitra, Q. Vo, and T. Tran (2021) Bouncing-to-wetting transition of water droplets impacting soft solids. Sodt Matter. Cited by: Conclusions & outlook.
  • M. A. Nearing, J. M. Bradford, and R. D. Holtz (1986) Measurement of force vs. time relations for waterdrop impact. Soil Sci. Soc. Am. J. 50, pp. 1532–1536. Cited by: §1.
  • M. J. Negus, M. R. Moore, J. M. Oliver, and R. Cimpeanu (2021) Droplet impact onto a spring-supported plate: analysis and simulations. J. Eng. Math. 128 (1), pp. 3. Cited by: §4.2.
  • M. Negus (2022) Modelling droplet impact in fluid-structure interaction problems. Ph.D. Thesis, University of Oxford. Cited by: §4.2.
  • J. G. Oldroyd and A. H. Wilson (1950) On the formulation of rheological equations of state. Proc. R. Soc. Lond. A 200 (1063), pp. 523–541. External Links: Document Cited by: §2.1.
  • Y.-H. Pao (1955) Extension of the Hertz theory of impact to the viscoelastic case. J. Appl. Phys. 26 (9), pp. 1083–1088. Cited by: §1, Conclusions & outlook.
  • J. Philippi, P.-Y. Lagrée, and A. Antkowiak (2016) Drop impact on a solid surface: short-time self-similarity. J. Fluid Mech. 795, pp. 96–135. Cited by: §1.
  • R. J. Poole (2012) The Deborah and Weissenberg numbers. Rheol. Bull. 53 (2), pp. 32–39. External Links: Link Cited by: §5.
  • S. Popinet and collaborators (2013) Basilisk C: volume of fluid method. Note: http://basilisk.fr (Last accessed: August 23, 2023) Cited by: §2.2.
  • S. Popinet (2018) Numerical models of surface tension. Annu. Rev. Fluid Mech. 50, pp. 49–75. Cited by: §2.2.
  • S. Popinet (2009) An accurate adaptive solver for surface-tension-driven interfacial flows. J. Comput. Phys. 228 (16), pp. 5838–5866. External Links: Document, ISSN 0021-9991, Link Cited by: §2.2, §2.2.
  • S. Popinet (2015) A quadtree-adaptive multigrid solver for the serre–green–naghdi equations. J. Comput. Phys. 302, pp. 336–358. External Links: Document, ISSN 0021-9991 Cited by: §2.2.
  • D. Richard and D. Quéré (2000) Bouncing water drops. Europhys. Lett. 50 (6), pp. 769–775. Cited by: Conclusions & outlook.
  • V. Sanjay, P. Chantelot, and D. Lohse (2023) When does an impacting drop stop bouncing?. J. Fluid Mech. 958, pp. A26. External Links: Document Cited by: Conclusions & outlook.
  • V. Sanjay, B. Zhang, C. Lv, and D. Lohse (2025) The role of viscosity on drop impact forces on non-wetting surfaces. J. Fluid Mech. 1004, pp. A6. External Links: Document Cited by: §1, §1, Fig. 2, Fig. 2, §3, §3, §4.2, §4.2, Conclusions & outlook.
  • V. Sanjay (2022) Viscous free-surface flows. PhD Thesis - Research UT, graduation UT, University of Twente, University of Twente, Netherlands. External Links: Document Cited by: §2.2, §3.
  • P. K. Sharma and H. N. Dixit (2021) Regimes of wettability-dependent and wettability-independent bouncing of a drop on a solid surface. J. Fluid Mech. 908, pp. A37. Cited by: Conclusions & outlook.
  • S. Shiri and J. C. Bird (2017) Heat exchange between a bouncing drop and a superhydrophobic substrate. Proc. Natl. Acad. Sci. USA 114 (27), pp. 6930–6935. External Links: Document Cited by: §1.
  • F. R. Smith, C. Nicloux, and D. Brutin (2018) Influence of the impact energy on the pattern of blood drip stains. Phys. Rev. Fluids 3, pp. 013601. External Links: Document, Link Cited by: §1.
  • F. T. Smith, L. Li, and G. X. Wu (2003) Air cushioning with a lubrication/inviscid balance. J. Fluid Mech. 482, pp. 291–318. Cited by: Conclusions & outlook, Conclusions & outlook.
  • J. H. Snoeijer and B. Andreotti (2013) Moving contact lines: scales, regimes, and dynamical transitions. Annu. Rev. Fluid Mech. 45, pp. 269–292. Cited by: Conclusions & outlook.
  • J. H. Snoeijer, A. Pandey, M. A. Herrada, and J. Eggers (2020) The relationship between viscoelasticity and elasticity. Proc. R. Soc. A 476 (2243), pp. 20200419. External Links: Document Cited by: §1, §2.1, §2.1, §2.1, Conclusions & outlook.
  • J. E. Sprittles (2024) Gas microfilms in droplet dynamics: when do drops bounce?. Annu. Rev. Fluid Mech. 56, pp. 91–118. Cited by: Conclusions & outlook, Conclusions & outlook.
  • S. Tamim and J. Bostwick (2021) Oscillations of a soft viscoelastic drop. npj Microgravity 7, pp. 42. External Links: Document Cited by: §1.
  • R.I. Tanner (2000) Engineering rheology. Oxford Engineering Science Series, OUP Oxford. External Links: ISBN 9780191590160, Link Cited by: §2.1, Conclusions & outlook.
  • N. P. Thien and R. I. Tanner (1977) A new constitutive equation derived from network theory. J. Non-Newton. Fluid 2 (4), pp. 353–365. External Links: Document, ISSN 0377-0257 Cited by: §2.1, Conclusions & outlook.
  • J. Thompson and J. Rothstein (2007) The atomization of viscoelastic fluids in flat-fan and hollow-cone spray nozzles. J. Non-Newtonian Fluid Mech. 147, pp. 11–22. External Links: Document Cited by: §1.
  • S. T. Thoroddsen, T. G. Etoh, K. Takehara, N. Ootsuka, and Y. Hatsuki (2005) The air bubble entrapped under a drop impacting on a solid surface. J. Fluid Mech. 545, pp. 203–212. External Links: Document Cited by: Conclusions & outlook.
  • S. T. Thoroddsen, T. G. Etoh, and K. Takehara (2003) Air entrapment under an impacting drop. J. Fluid Mech. 478, pp. 125–134. Cited by: Conclusions & outlook.
  • Y. Tirupataiah and G. Sundararajan (1991) A dynamic indentation technique for the characterization of the high strain rate plastic flow behaviour of ductile metals and alloys. J. Mech. Phys. Solids 39 (2), pp. 243–271. External Links: Document, ISSN 0022-5096 Cited by: §1.
  • G. Tryggvason, R. Scardovelli, and S. Zaleski (2011) Direct Numerical Simulations of Gas–Liquid Multiphase Flows. Cambridge University Press. Cited by: §2.2.
  • R. C. A. van der Veen, T. Tran, D. Lohse, and C. Sun (2012) Direct measurements of air layer profiles under impacting droplets using high-speed color interferometry. Phys. Rev. E 85, pp. 026315. Cited by: Conclusions & outlook, Conclusions & outlook.
  • H. Wagner (1932) Über Stoß- und Gleitvorgänge an der Oberfläche von Flüssigkeiten. Z. Angew. Math. Mech. 12 (4), pp. 193–215. Cited by: §1.
  • B. Zhang, J. Li, P. Guo, and Q. Lv (2017) Experimental studies on the effect of reynolds and weber numbers on the impact forces of low-speed droplets colliding with a solid surface. Exp. Fluids 58 (9), pp. 1–12. Cited by: Conclusions & outlook.
  • B. Zhang, V. Sanjay, S. Shi, Y. Zhao, C. Lv, and D. Lohse (2022) Impact forces of water drops falling on superhydrophobic surfaces. Phys. Rev. Lett. 129, pp. 104501. Cited by: §1, §1, Fig. 2, Fig. 2, §3, §3, §3, §4.2, §4.2, Conclusions & outlook.
  • L. Zhang, T. Soori, A. Rokoni, A. Kaminski, and Y. Sun (2021) Thin film instability driven dimple mode of air film failure during drop impact on smooth surfaces. Phys. Rev. Fluids 6, pp. 044002. Cited by: Conclusions & outlook.
  • S. Zheng, S. Dillavou, and J. M. Kolinski (2021) Air mediates the impact of a compliant hemisphere on a rigid smooth surface. Soft Matter 17 (14), pp. 3813–3819. Cited by: §2.2, Conclusions & outlook.